OpenStudy (anonymous):

need help plz Evaluate the following: 3^2 ÷ (2 + 1). 2 3 4 5

OpenStudy (anonymous):

5 maybe?

OpenStudy (anonymous):

are you sure and why plz

OpenStudy (koikkara):

\(3^2 ÷ (2 + 1)\)= \(9 ÷ 3\) = \(?\)

OpenStudy (koikkara):

@itsnotjesse

OpenStudy (anonymous):

9/3= 3

OpenStudy (anonymous):

using a calculator i came to 3

OpenStudy (koikkara):

\(\color{blue}{\huge\tt{Welcome~}}\) \(\huge\mathcal{to}\) \(\color{red}{\huge\tt{Open~Study~!}}\) @itsnotjesse Good !! \(\color{blue}{\huge\tt{Nice~}}\) \(\huge\mathcal{to~meet~}\) \(\color{red}{\huge\tt{You~!}}\)

OpenStudy (anonymous):

oh ok thank you that was easier than i thought

OpenStudy (anonymous):

3^2= 9/3=3 (2+1=3)^

OpenStudy (anonymous):

@Koikkara can you help me with one more

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

ok thank you one sec

OpenStudy (anonymous):

4(a ⋅ b) = (4 ⋅ a)b is an example of which algebraic property? Associative Property of Multiplication Transitive Property Reflexive Property Commutative Property of Multiplication i think its a or d

OpenStudy (anonymous):

it is the commutative property of multiplication :)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

no problem!